OpenStudy (anonymous):

okay, so if a graph has the equation sin(bx)=0.5 and they show me what the graph looks like, how would I go about finding 'b'? I really have no idea.

OpenStudy (anonymous):

this is what the graph looks like

OpenStudy (anonymous):

oooh i see it is the graph of \[y=\sin(bx)-.5\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok you had =.5 and that confused me you need to find the period of this fuction. find it from the graph

OpenStudy (anonymous):

and that will give me b?

OpenStudy (anonymous):

no there is anohter step first. it looks like the period is 180

OpenStudy (anonymous):

in general if you have \(y=\sin(bx)\) the period is \(\frac{2\pi}{b}\) but it looks like you are working in degrees so the period would be \(\frac{360}{b}\)

OpenStudy (anonymous):

so since the period is 180, set \(\frac{360}{b}=180\) solve for b, get \(b=2\)

OpenStudy (anonymous):

you can see the graph of \(y=\sin(2x)-.5\) here. looks like what you got http://www.wolframalpha.com/input/?i=sin%282x%29-.5

OpenStudy (anonymous):

THANK YOU! that actually makes so much sense...i thought that since it was a multiple angle question there would be something else that you would need to do to get the answer!

OpenStudy (anonymous):

yw